If you could show work it would be greatly appreciated

SallyMay, Inc., designs and manufactures T-shirts. It sells its T-shirts to brand name clothes retailers in lots of one dozen. SallyMay's May 2013 static budget and actual results for direct inputs are as follows:

Static Budget
Number of T-shirt lots (1 lot1 dozen) 400

Per Lot of T-shirts:
Direct materials 14 meters at $1.70 per meter$23.80
Direct manufacturing labor 1.6 hours at $8.10 per hour $12.96

Actual Results
Number of T-shirt lots sold 450

Total Direct Inputs:
Direct materials 6,840 meters at $1.95 per meter = $13,338
Direct manufacturing labor 675 hours at $8.20 per hour = $5,535

SallyMay has a policy of analyzing all input variances when they add up to more than 10% of the total cost of materials and labor in the flexible budget, and this is true in May 2013. The production manager discusses the sources of the variances: "A new type of material was purchased in May. This led to faster cutting and sewing, but the workers used more material than usual as they learned to work with it. For now, the standards are fine."

Calculate the direct materials and direct manufacturing labor price and efficiency variances in May 2013. What is the total flexible-budget variance for both inputs (direct materials and direct manufacturing labor) combined? What percentage is this variance of the total cost of direct materials and direct manufacturing labor in the flexible budget?

Sally King, the CEO, is concerned about the input variances. But she likes the quality and feel of the new material and agrees to use it for one more year. In May 2014, SallyMay again produces 450 lots of T-shirts. Relative to May 2013, 2% less direct material is used, direct material price is down 5%, and 2% less direct manufacturing labor is used. Labor price has remained the same as in May 2013. Calculate the direct materials and direct manufacturing labor price and efficiency variances in May 2014. What is the total flexible-budget variance for both inputs (direct materials and direct manufacturing labor) combined? What percentage is this variance of the total cost of direct materials and direct manufacturing labor in the flexible budget?

Comment on the May 2014 results. Would you continue the "experiment" of using the new material?

4-variance analysis, fill in the blanks.
Variable U/F Fixed U/F
1. Spending variance
2. Efficiency variance
3. Production-volume variance
4. Flexible-budget variance
5. Underallocated (overallocated) MOH
You can use the worksheet below to calculate the variances for the above table
Flexible Budget: Allocated:
Budgeted Input Qty. Budgeted Input Qty.
Allowed for Allowed for
Actual Costs Actual Input Qty. Actual Output Actual Output
Incurred × Budgeted Rate × Budgeted Rate × Budgeted Rate
Variable Manufacturing Overhead
Spending variance
Efficiency Variance
Never a Variance
Flexible-budget variance
Never a Variance
Overallocated Variable Overhead
Flexible Budget:
Same Budgeted Same Budgeted Allocated:
Lump Sum Lump Sum Budgeted Input Qty.
(as in Static Budget) (as in Static Budget) Allowed for
Regardless of Regardless of Actual Output
Actual Costs Incurred Output Level Output Level × Budgeted Rate
Fixed Manufacturing Overhead
Spending variance
Never a Variance
Production volume Variance
Flexible-budget variance
Production volume Variance
Underallocated fixed Overhead
Answer #1

For May 2013

Static Budget 400 Lots

Direct Material = 14 meter @ 1.70 per meter

Direct Labour = 1.6 Hours @8.10 per hour

Actual result 450 lots

Direct Material = 6840 meter @ 1.95 per meter

Direct Labour = 675 hours @ 8.20 per hours

Direct Material Price Variance = Actual Quantity (Standard Rate - Actual Rate)

= 6840 (1.70 - 1.95)

= 1,710 (U)

Direct Material usage Variance = Standard Rate (Standard Quantity - Actual Quantity)

= 1.70 (14*450 - 6840)

= 918 (U)

Direct Labour Rate Variance = Actual Hour (Standard Rate - Actual Rate)

= 675 (8.10 - 8.20)

= 67.5 (U)

Direct Labour Efficiency Variance = Standar Rate (Standard Hour - Actual Hour)

= 8.10 (640 - 675)

= 283.5 (U)

Total Flexible budget for Material = 1710 + 918 = 2,628 (U)

Total Flexible budget for Labour = 67.5 + 283.5 = 351

In percentage:

Material Variance = 2628 * 100 / 13338 = 19.70%

Labour Variance = 351 * 100 / 5535 = 6.34%

For May 2013

Actual result 450 lots

Direct Material = 6703.2 meter @ 1.85 per meter

Direct Labour = 661.5 hours @ 8.20 per hours

Direct Material Price Variance = Actual Quantity (Standard Rate - Actual Rate)

= 6703.2 (1.70 - 1.85)

= 1005.48 (U)

Direct Material usage Variance = Standard Rate (Standard Quantity - Actual Quantity)

= 1.70 (14*450 - 6703.2)

= 685.44 (U)

Direct Labour Rate Variance = Actual Hour (Standard Rate - Actual Rate)

= 661.5 (8.10 - 8.20)

= 66.15 (U)

Direct Labour Efficiency Variance = Standar Rate (Standard Hour - Actual Hour)

= 8.10 (640 - 661.5)

= 174.15 (U)

Total Flexible budget for Material = 1005.48 + 685.44 = 1690.92 (U)

Total Flexible budget for Labour = 66.15 + 174.15 = 240.3

In percentage:

Material Variance = 1690.92 * 100 / 13338 = 12.68%

Labour Variance = 240.3 * 100 / 5535 = 4.34%

As per may 2014 result, SallyMay should continue the "experiment" of using new Material.